OpenStudy (anonymous):

Solve by completing the square & quadratic formula: x^2 6x + 17 = 0

OpenStudy (anonymous):

\[x^2+6x+17=0\] or \[x^2-6x+17=0\]??

OpenStudy (anonymous):

Gah, sorry. x^2 + 6x + 17 = 0

OpenStudy (anonymous):

\[x^2+6x=-17\] \[(x+3)^2=-17+9\] \[(x+3)^2=-8\] now either you are done because you can't take the root of a negative number, or else you are allowing complex solutions. which is it?

OpenStudy (anonymous):

I'm assuming the second option- some of the answers are things like: \[x = 3i \sqrt{2}\]

OpenStudy (anonymous):

actually if you are using complext numbers it would be \[x=2\sqrt{2}i\] or \[x=-2\sqrt{2}i\]

OpenStudy (anonymous):

...What is it with me today...? But, yes, we've been doing complex numbers.